### v1

# A Two-Stage Pretraining-Finetuning Framework for Treatment Effect Estimation with Unmeasured Confounding

Abstract.
Estimating the conditional average treatment effect (CATE) from observational data plays a crucial role in areas such as e-commerce, healthcare, and economics. Existing studies mainly rely on the strong ignorability assumption that there are no unmeasured confounders, whose presence cannot be tested from observational data and can invalidate any causal conclusion. In contrast, data collected from randomized controlled trials (RCT) do not suffer from confounding, but are usually limited by a small sample size. In this paper, we propose a two-stage pretraining-finetuning (TSPF) framework using both large-scale observational data and small-scale RCT data to estimate the CATE in the presence of unmeasured confounding. In the first stage, a foundational representation of covariates is trained to estimate counterfactual outcomes through large-scale observational data. In the second stage, we propose to train an augmented representation of the covariates, which is concatenated to the foundational representation obtained in the first stage to adjust for the unmeasured confounding. To avoid overfitting caused by the small-scale RCT data in the second stage, we further propose a partial parameter initialization approach, rather than training a separate network. The superiority of our approach is validated on two public datasets with extensive experiments. The code is available at https://github.com/zhouchuanCN/KDD25-TSPF.
Causal Effect Estimation, Unmeasured Confounding, Data Fusion
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709161isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Machine learning approaches
1. Introduction
The conditional average treatment effect (CATE) is the average causal effect of a treatment or an intervention on the outcome of interest given the covariates (Pearl, 2009), which plays an important role in diverse fields, such as e-commerce (Wu et al., 2022b), healthcare (Robins and Hernán, 2016), and economics (Huynh et al., 2016). In e-commerce, the platforms desire to predict how recommending a specific product to a particular user affects the probability of purchase (Li et al., 2022), and thereby influence the total profit. In healthcare, doctors assess the potential outcome for different patient groups when administering a certain treatment (Corrao et al., 2012) for precision medicine. Similarly in economics, the policymakers evaluate how much a job training program will raise employment opportunities for unemployed individuals (Britto et al., 2022).
To enhance the accuracy of CATE estimation, representation-based learning approaches have gathered increasing attention due to their impressive performance (Johansson et al., 2016; Li et al., 2024a; Zhu et al., 2024a; Wang et al., 2024). These approaches focus on generating covariate representations, with the objective of mitigating confounding bias by minimizing distributional discrepancies between the treatment group and control group. To obtain such representations, previous approaches have developed substantial theory and explored extensive practice. For instance, some of them use integral probability metric (IPM) for regularization (Johansson et al., 2016), while a few approaches emphasize local similarity preservation (Yao et al., 2018), targeted learning (Zhang et al., 2020), and optimal transport (Wang et al., 2024).
However, the aforementioned methods may ignore unmeasured confounding, which is very common in real-world scenarios. In our e-commerce example, the financial status of users should be sensitive and cannot be collected (Li et al., 2023). In the healthcare case, the personal lifestyles of patients are difficult to obtain (Charpignon et al., 2022). For the example in economics, personal working status is difficult to measure (Xu and Li, 2020). These unmeasured variables can affect the treatment and outcome simultaneously, which causes confounding bias. Therefore, proposing methods to account for the confounding bias is crucial to accurately estimate CATE.
To address unmeasured confounding, one category of mainstream methods can only rely on large-scale observational (OBS) data, including sensitivity analysis (Imbens, 2003; Dorn et al., 2024), front door adjustment methods (Fulcher et al., 2020; Zhu et al., 2024b), and instrumental variables methods (Angrist et al., 1996; Rudolph et al., 2024). These methods require additional strong assumptions that cannot be tested from the data and raise concerns if these assumptions are violated (Hartwig et al., 2023; Kong et al., 2025). Compared to the OBS data, randomized controlled trial (RCT) data are considered as the gold standard for causal effect estimation (Prosperi et al., 2020). However, practical challenges such as high costs and ethical concerns may make the collection of RCT data difficult (Zabor et al., 2020; Bédécarrats et al., 2020), resulting in limited sample sizes. Due to the small size, it is impractical to directly train causal effect prediction models on RCT data alone (Hoogland et al., 2021). Therefore, it is necessary to find an effective method to combine small-scale RCT data with large-scale OBS data. Some studies use methods of correcting for residuals, in which RCT data are utilized to correct for biased estimates from OBS data (Colnet et al., 2024). However, these methods still rely on strong assumptions, including linear and additive data generation processes (Kallus et al., 2018) and a shared structure between the two data types (RCT and OBS) (Hatt et al., 2022b). This can lead to poor model performance in the complex real world. In contrast, other studies without these assumptions generally use both types of data to estimate CATE and the residuals between the biased and unbiased estimates (Wu and Yang, 2022), yet these methods suffer from the risk of overfitting due to insufficient RCT sample sizes.
In this paper, we introduce a two-stage framework named TSPF for CATE estimation with unmeasured confounding based on the pretraining-finetuning principle. Our approach leverages large-scale OBS data to train a foundational representation of covariates and then uses relatively small-scale RCT data to adjust the representation learned from OBS data. We then train a more accurate prediction model using this adjusted representation. In the second stage, we introduce an additional module that ensures stronger representation ability compared to the methods that use RCT data to estimate the residuals between the biased and unbiased estimates.
The contributions of this paper are summarized as follows.
  • We present a two-stage pretraining-finetuning framework for estimating the CATE. This framework tackles the issue of unmeasured confounding by using a small amount of unconfounded RCT data to calibrate the representations learned from observational (OBS) data.
  • The proposed framework does not rely on the linear and additive generation assumptions, and can flexibly adjust its model structure according to the sample size of RCT data, thus mitigating the over-fitting problem.
  • Extensive experiments conducted on the IHDP and Jobs datasets demonstrate the effectiveness of our approach.
2. Related Work
2.1. CATE Estimation
CATE also known as heterogeneous treatment effect (HTE), refers to the average causal effects of a treatment on an outcome for subgroups with different covariates. To accurately estimate CATE, early statistical methods include matching (Dehejia and Wahba, 2002), stratification (O’Muircheartaigh and Hedges, 2014), reweighting (Rosenbaum, 1987; Bang and Robins, 2005), and tree-based methods (Chipman et al., 2010; Wager and Athey, 2018). With the impressive advances in deep learning, recent CATE estimation methods can be broadly categorized into representational learning-based and generative model-based methods. Specifically, the core idea of representation learning-based methods is to learn a balanced covariate representation that has similar distributions in the treatment and control groups to mitigate confounding bias (Assaad et al., 2021; Yao et al., 2021; Guo et al., 2020). Such balanced representations can be obtained by using integral probability metric (IPM) for regularization (Johansson et al., 2016; Shalit et al., 2017), local similarity preservation (Yao et al., 2018, 2019), targeted learning (Shi et al., 2019; Zhang et al., 2020), and optimal transport (Wang et al., 2024; Torous et al., 2021). An alternative category is generative model-based methods that estimate the counterfactual outcomes by assuming their data generation process and exploiting generative models (Zhang et al., 2021; Zou et al., 2020; Guo et al., 2020). Specifically, CEVAE uses variational autoencoder (VAE) to learn confounders from observed variables (Louizos et al., 2017). SCIGAN generates missing counterfactual results based on generative adversarial networks (GAN) and combines the facts and estimated counterfactual results to estimate CATE (Bica et al., 2020). Unlike previous representation learning-based methods, we do not require no unmeasured confounding assumption. Instead, we learn an augmented representation to address the effects of unmeasured confounding for correcting biased estimates.
2.2. Unmeasured Confounding
Unmeasured confounding refers to a situation where there are unmeasured variables in the study that influence both the treatment and the outcome, which may lead to bias in CATE estimates (Ananth and Schisterman, 2018). To address this problem, previous studies can be broadly categorized into two types. The first only uses OBS data and mainly includes sensitivity analysis and auxiliary information methods. Sensitivity analysis aims to quantify the potential impact of unmeasured confounding on the treatment effect and to obtain a bound on the treatment effect (Rosenbaum and Rubin, 1983; Robins et al., 2000). However, these approaches assume a determined confounding mechanism of the unmeasured variables (Imbens, 2003). This assumption is too strong and can easily be violated in practice (Franks et al., 2019; Veitch and Zaveri, 2020). Auxiliary information methods in causal inference mainly include instrumental variable (IV) methods and front-door adjustment methods (Li et al., 2024c). IV methods rely on external variables to address unmeasured confounding in observational studies (Imbens, 2014; Wu et al., 2022a; Rudolph et al., 2024). However, these methods assume linearity and require unconfounded instruments, posing practical limitations (Frauen and Feuerriegel, 2022). Front door adjustment methods estimate the causal effect of a treatment on an outcome by leveraging a causal pathway, also known as the front-door criterion, which blocks all possible influences of unmeasured confounders (Bellemare et al., 2020; Fulcher et al., 2020). In general, front door adjustment may rely on knowing the true causal graph, which may not always be feasible in practice (Shah et al., 2024; Li et al., 2024b). Another type of approaches combine both OBS and RCT data to mitigate unmeasured confounding. Kallus et al. (Kallus et al., 2018) propose to use RCT data to correct the bias in CATE estimates derived from OBS data under the assumption that the unmeasured confounding bias is a linear function of the covariates (Kallus et al., 2018). Hatt et al. (Hatt et al., 2022a) extend to the non-linear case by using shared model structure between OBS and RCT data. Rather than using OBS and RCT data in separate regressions, by joint modeling the unmeasured confounding bias and the CATE with OBS and RCT data simultaneously as input, Wu et al. (Wu and Yang, 2022) identify both the bias from OBS estimates and unbiased CATE estimates using the R-learner approach, and Cheng et al. (Cheng and Cai, 2021) perform a weighted average of the estimates from the two data sources to mitigate the bias. Both the determination of weights and the estimation of CATE require a large sample size. However, this requirement is at odds with the small sample size of RCT. In this paper, we adopt a more flexible network architecture, with a deep neural network for predicting biased outcomes using the OBS data and extra neural networks for adjusting the representation as well as the prediction using the RCT data, which allows us to utilize limited RCT data to achieve better performance.
3. Preliminaries
3.1. Problem Setup
We consider two independent data sources taken from the same target population: one from OBS and the other from RCT. Each individual in the OBS or RCT study is an observation of $(X,Y,T,G)$, a random tuple with distribution $P$. For the $i$-th individual, the observation comprises $d$-dimensional covariates $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, the observed outcome $Y_{i}$, the assigned binary treatments $T_{i}\in\{0,1\}$ ($T_{i}=0$ for the controlled and $T_{i}=1$ for the treated individuals) and $G_{i}$ denoting participation in the OBS ($G_{i}=0$) or RCT ($G_{i}=1$) study. Using the Neyman-Rubin potential outcome framework (Imbens and Rubin, 2015), we let $Y^{1}_{i},Y^{0}_{i}$ be the potential outcomes. We denote the OBS data as $\mathcal{D}^{OBS}=\{(X_{i},T_{i},Y_{i},G_{i}=0):i\in\mathcal{O}\}$ with sample size $n$, and the RCT data as $\mathcal{D}^{RCT}=\{(X_{i},T_{i},Y_{i},G_{i}=1):i\in\mathcal{R}\}$ with sample size $m$, where $\mathcal{O}=\{1,\ldots,n\}$ and =$\mathcal{R}=\{n+1,\ldots,n+m\}$ are sample index sets for the OBS and RCT data, respectively. The total sample size is $N=n+m$. We define the propensity score as $e(x,G)=P(T=1\mid X=x,G)$. The CATE is defined as the conditional expectation of difference between potential outcomes under the treatment group and the control group as follows:
$\tau(x)=\mathbb{E}[Y^{1}-Y^{0}\mid X=x].$
3.2. Identification of CATE
To identify the CATE from observed data, in addition to the Stable Unit Treatment Value Assumption (SUTVA) that there are no interference between units and there are no different forms of each treatment level, the following three assumptions are required:
Assumption 1.
(Ignorability) $(Y^{1},Y^{0})\perp\!\!\!\perp T\mid X$,
Assumption 2.
(Consistency) $Y=TY^{1}+(1-T)Y^{0}$,
Assumption 3.
(Positivity) $10<e(x,G)<$ 1 for all $x\in\mathcal{X}$.
Assumption 1 is also known as no unmeasured confounding, which holds in the RCT by default due to the randomized treatment assignment. We can identify CATE based on RCT:
$\tau(x)=\mathbb{E}[Y\mid T=1,X=x,G=1]-\mathbb{E}[Y\mid T=0,X=x,G=1].$
[…]
$\omega(x)=\mathbb{E}\left[Y\mid T=1,X=x,G=0\right]-\mathbb{E}\left[Y\mid T=0,X=x,G=0\right].$
Note that due to unmeasured confounding, $\omega(x)\neq\tau(x)$ for any $x$. The difference between these two quantities is precisely the confounding effect, which we denote the residual function as:
$\eta(x)=\tau(x)-\omega(x).$
In each treatment group, we denote the residual functions:
$\eta_{t}(x)=\mathbb{E}[Y\mid T=1,X=x,G=0]-\mathbb{E}[Y\mid T=1,X=x,G=1],$
$\eta_{c}(x)=\mathbb{E}[Y\mid T=0,X=x,G=0]-\mathbb{E}[Y\mid T=0,X=x,G=1].$
3.3. Previous Work
To address the unmeasured confounding, previous work proposes many methods to estimate the residual function $\eta(x)$ and obtain an accurate estimation of CATE $\tau(x)$ with extra small-scale RCT data.
3.3.1. Residual Correction
This type of approaches first only uses the OBS data for initial estimation and then uses RCT data for correcting the biased estimation with a residual function (Kallus et al., 2018).
Step 1: Estimate $\omega(x)$ using the OBS data, denoted as $\hat{\omega}(x)$.
Step 2: Estimate $\eta(x)$ using the RCT data by minimizing:
$\min_{\eta}\sum_{i\in\mathcal{R}}(Y_{i}^{*}-\hat{\omega}(X_{i})-\eta(X_{i}))^{2},$
where $Y^{*}_{i}=\frac{T_{i}Y_{i}}{e(X_{i},G_{i})}-\frac{(1-T_{i})Y_{i}}{1-e(X_{i},G_{i})}$ is the pseudo-outcome.
The final estimate of $\tau(x)$ is $\hat{\omega}(x)+\hat{\eta}(x)$. To ensure a consistent estimate of $\tau(x)$, the residual $\eta(x)$ is assumed to be linear and additive. Moreover, the small-scale RCT data used in the second phase may cause overfitting of the estimated residual $\hat{\eta}(x)$.
3.3.2. Joint Learning of Residual and CATE Models
This type of approaches involves training the models for estimating $\eta(x)$ and $\tau(x)$ simultaneously using both OBS and RCT data (Wu and Yang, 2022).
Step 1: Estimate the conditional average outcome $\mu(X,G)$ and propensity score $e(X,G)$ by using the combined OBS and RCT data, denoted as $\hat{\mu}(X,G)$ and $\hat{e}(X,G)$.
Step 2: Estimate $\{\eta(\cdot),\tau(\cdot)\}$ by the optimization problem:
$\arg\min_{\eta,\tau}\sum_{i\in\mathcal{O}\cup\mathcal{R}}$$\left\{Y_{i}-\hat{\mu}\left(X_{i},G_{i}\right)-[\tau(X_{i})+(1-G_{i})\eta(X_{i})]\right.$
$\cdot\left.[T_{i}-\hat{e}(X_{i},G_{i})]\right\}^{2}+\Lambda(\tau)+\Lambda(\eta),$
where $\Lambda(\tau)$ and $\Lambda(\eta)$ are regularization terms on the complexity of the $\tau(\cdot)$ and $\eta(\cdot)$ functions.
Since RCT generally involves much smaller sample sizes compared to OBS studies, the data available for explicitly learning $\eta(x)$ is limited. Insufficient RCT data can lead to overfitting and a high variance of $\hat{\eta}(x)$, thus reducing the accuracy of CATE estimation.
Figure 1. The framework of our proposed method, which is composed of the modules for the first stage (blue) and second stage (yellow). Note that the two $\phi^{U}$ shown in the figure represent the same module.
4. Methodology
In this section, we present a two-stage framework for CATE estimation based on the pretraining-finetuning principle, as shown in Figure 1. The motivation is to use large-scale OBS data to train a base representation of covariates, then use relatively small-scale unbiased RCT data to calibrate the bias in the representation learned from OBS data for training the unbiased prediction model. Specifically, in the first stage, only the OBS data is used. We start with a representation module $\phi$, followed by one reconstruction module $\psi$ and two prediction heads $h_{0}$ and $h_{1}$ to ensure the learned covariate representation can have enough information and predict the outcome for control group and treatment group simultaneously. While in the second stage, we use only the RCT data. The representation module $\phi$ learned in the first stage is frozen, with a learnable representation adapter module $\phi^{U}$ to calibrate the bias of $\phi$. Then the concatenated representation is fed to two prediction heads $g_{0}$ and $g_{1}$ to obtain the unbiased predicted potential outcomes under control and treatment groups respectively. We carefully design an initialization strategy to ensure that the initialized second-stage model produces the same predictions as the converged model in the first stage. Our approach distinguishes from the one proposed by Kallus et al. (Kallus et al., 2018), which only uses linear regression to estimate the residual function $\eta(x)$ in the second stage. We can regard the first stage as pretraining on the large OBS data and the second stage as finetuning on the small-scale unbiased RCT data.
4.1. First Stage: Pretraining Stage
The goal of our first-stage training is to obtain a representation module as well as prediction heads that can accurately estimate the potential outcomes of OBS data. These modules offer high-quality initialization for second-stage training, allowing fine-tuning on the RCT data to avoid the overfitting problem. A three-headed architecture and a multi-task training framework are employed to achieve this goal. Next, we will look into the details of each module.
4.1.1. Representation
We design a multi-layer feed-forward neural network $\phi$ to obtain a representation $Z$ for the covariates $X$ for both treatment and control groups. In other words, for an individual sample $(X_{i},T_{i},Y_{i},G_{i}=0)$, the representation $Z_{i}=\phi(X_{i})$ remains the same whether $T_{i}=0$ or $T_{i}=1$. To better estimate the causal effect, we adopt the idea of covariate balancing on the representation space as Shalit et al. (Shalit et al., 2017) Specifically, the representations for treatment group $\{Z_{i}=\phi(X_{i}):G_{i}=0,T_{i}=1\}$ are regarded as i.i.d samples drawn randomly from a distribution $P_{\phi}^{t=1}$ and similarly $P_{\phi}^{t=0}$ for the control group. We anticipate the distributions of representations to be similar between the treatment and control groups. An integral probability metric (IPM) is employed to measure the distance between the two distributions. Thus the covariate imbalancing loss is defined as:
$\mathcal{L}_{imb}=\mathrm{IPM}_{\mathcal{G}}(\hat{P}_{\phi}^{t=0},\hat{P}_{\phi}^{t=1}),$
where $\mathrm{IPM}_{\mathcal{G}}(\cdot)$ is the empirical IPM defined by the function family $\mathcal{G}$, and $\hat{P}_{\phi}^{t=1}$ and $\hat{P}_{\phi}^{t=0}$ are empirical distributions of $P_{\phi}^{t=1}$ and $P_{\phi}^{t=0}$ respectively. In the implementation, we adopt Wasserstein distance as a showcase, which can be consistently estimated from finite samples within a mini-batch (Frogner et al., 2015).
4.1.2. Reconstruction and Prediction
To ensure that $Z$ retains as much information about the original covariates as possible, we introduce the decoder network $\psi$ to reconstruct the original covariates: $\hat{X}=\psi(Z)=\psi(\phi(X))$. The reconstruction loss is computed by the mean squared error (MSE):
$\mathcal{L}_{rec}=\dfrac{1}{|\mathcal{O}|}\sum_{i\in\mathcal{O}}||\hat{X}_{i}-X_{i}||_{2}^{2},$
where $\hat{X}_{i}=\psi(\phi(X_{i}))$ is reconstructed covariate for the $i$-th sample in the OBS data. The reconstruction design resembling an autoencoder allows the learned representations to encompass nearly complete information in the covariates, rather than only the information necessary for fitting the training set, thereby enhancing the generalization of our representation module.
We then use the representations $Z$ to estimate the potential outcomes with two $l_{p}-$layer prediction heads $h_{0}$ and $h_{1}$, which are the predictors for control and treatment outcomes, respectively. Note that unmeasured confounding in the observational data can lead to biased estimations of potential outcomes, we refer to the prediction result $\tilde{Y}^{0}=h_{0}(Z)=h_{0}(\phi(X))$ as the pseudo control outcome and $\tilde{Y}^{1}=h_{1}(Z)=h_{1}(\phi(X))$ as the pseudo treatment outcome. To enhance comparability between the treatment group and control group, we employ a reweighting technique to balance the two groups. Formally, let $f_{h}(x,t)=h_{t}(\phi(x))$ with $t\in\{0,1\}$ be the predicted potential outcomes by via the two heads $h_{0}$ and $h_{1}$, the loss for outcome prediction is as follows:
$\mathcal{L}_{f}=\frac{1}{|\mathcal{O}|}\sum_{i\in\mathcal{O}}w_{i}\cdot l(Y_{i},f_{h}(X_{i},T_{i})),$
with $w_{i}=\frac{T_{i}}{2u}+\frac{1-T_{i}}{2(1-u)}$, where $u=\frac{1}{n}\sum_{i=1}^{n}T_{i}$. The loss function $l(\cdot,\cdot)$ in $\mathcal{L}_{f}$ is flexible and can be determined based on the value range of potential outcomes. If the potential outcomes are binary, a cross-entropy loss is appropriate, whereas for continuous potential outcomes, an MSE loss is preferable.
In summary, in the first-stage training, we use the following training objective:
$\min_{\phi,\psi,h_{0},h_{1}}\mathcal{L}_{f}+\lambda_{1}\mathcal{L}_{rec}+\lambda_{2}\mathcal{L}_{imb},$
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are tunable hyperparameters.
4.2. Second Stage: Finetuning Stage
In the second stage of training, we exploit the small-scale unconfounded RCT data to remove the hidden confounding by concatenating the learned covariate representation in the first stage with a newly learned augmented covariate representation, then finetuning the prediction heads to obtain an unbiased CATE estimation. To achieve this, we keep the biased representation $Z$ produced by $\phi$ unchanged but only treat it as a part of the representation, together with an additional $Z^{U}$ generated by another representation module $\phi^{U}$. We call $\phi^{U}$ a representation adapter, as it helps to adapt the final representation to account for the hidden confounding in the observed data. In addition, a large proportion of the parameters of the prediction heads $g_{0}$ and $g_{1}$ are initialized by $h_{0}$ and $h_{1}$, respectively. With the above steps, the second stage aims at adjusting the hidden confounding through the augmented covariate representation and the finetuned prediction heads with partial parameter initialization. Below we explain the modules in detail.
4.2.1. Representation Adapter
We employ a shallower feed-forward network $\phi^{U}$ as the representation adapter. It is worth noting that the width and depth of $\phi^{U}$ can be adjusted based on the scale of RCT data size. If the size of RCT data was comparable to that of OBS data, we can use the same architecture as $\phi$. However, in real-world cases, the RCT data is rare compared to the OBS data, so the size of $\phi^{U}$ should be smaller. We denote the representation generated by $\phi^{U}$ as $Z^{U}$. To make sure that $Z^{U}$ captures different features of covariates from $Z$, we employ mutual information to control the overlap between the two covariate representations:
$\mathcal{L}_{MI}=CLUB(Z,Z^{U}),$
$CLUB(Z,Z^{U})$ $=\frac{1}{m^{2}}\sum_{i\in\mathcal{R}}\sum_{j\in\mathcal{R}}\Big{[}\log q_{\theta}(Z^{U}_{i}\mid Z_{i})-\log q_{\theta}(Z^{U}_{j}\mid Z_{i})\Big{]}$
$=\frac{1}{m}\sum_{i\in\mathcal{R}}\Big{[}\log q_{\theta}(Z^{U}_{i}\mid Z_{i})-\frac{1}{m}\sum_{j\in\mathcal{R}}\log q_{\theta}(Z^{U}_{j}\mid Z_{i})\Big{]},$
where $CLUB(Z,Z^{U})$ is the empirical Contrastive Log-ratio Upper Bound (CLUB) of mutual information (Cheng et al., 2020) between two covariate representations $Z$ and $Z^{U}$, and $q_{\theta}(Z^{U}\mid Z)$ is the variational approximation of $P(Z^{U}\mid Z)$. With good variational approximation $q_{\theta}(Z^{U}\mid Z)$, it can be shown that the empirical CLUB is still a valid upper bound of the ground-truth mutual information.
4.2.2. Prediction
Similarly to the first stage, we design two $l_{p}-$layer prediction heads $g_{0}$ and $g_{1}$ to estimate the potential outcomes under control and treatment groups with unmeasured confoundings, respectively. Notice that $g_{0},g_{1}$ and $h_{0},h_{1}$ have the same depth $l_{p}$, yet every layer of $g_{t}$ has a larger or equal width than $h_{t}$. For $t\in\{0,1\}$, we define the header $h_{t}$ as:
$a_{h_{t}}^{(0)}=Z,~{}a_{h_{t}}^{(l)}=\sigma(W_{h_{t}}^{(l)}a_{h_{t}}^{(l-1)}+b_{h_{t}}^{(l)}),\text{ for }l=1,2,\ldots,l_{p}-1,$
$a_{h_{t}}^{(l_{p})}=\tilde{Y}_{t}=W_{h_{t}}^{(l_{p})}a_{h_{t}}^{(l_{p}-1)}+b_{h_{t}}^{(l_{p})},$
where $W_{h_{t}}^{(l)}$ is the weight matrix from layer $l-1$ to layer $l$, $b_{h_{t}}^{(l)}$ is the bias vector of layer $l$, $a_{h_{t}}^{(l)}$ is the output of layer $l$ for $l\in\{1,2,\ldots,l_{p}\}$ and $\sigma$ is the activation function. The definition of $g_{t}$ is:
$a_{g_{t}}^{(0)}=\begin{bmatrix}Z\\ Z^{U}\end{bmatrix},~{}a_{g_{t}}^{(l)}=\sigma(W_{g_{t}}^{(l)}a_{g_{t}}^{(l-1)}+b_{g_{t}}^{(l)}),\text{ for }l=1,2,\ldots,l_{p}-1,$
$a_{g_{t}}^{(l_{p})}=\hat{Y}_{t}=W_{g_{t}}^{(l_{p})}a_{g_{t}}^{(l_{p}-1)}+b_{g_{t}}^{(l_{p})},$
with $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ having similar meanings as $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$ but for $g_{t}$, $\hat{Y}_{t}$ is the final prediction for potential outcome under treatment $t$. In our design, every layer of $g_{t}$ has a larger or equal width than $h_{t}$, thus the dimension of $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ is no less than that of $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$ respectively. We divide the parameters of $g_{t}$ as:
$W_{g_{t}}^{(l)}=\begin{bmatrix}{}^{1}W_{g_{t}}^{(l)},^{2}W_{g_{t}}^{(l)}\\ {}^{3}W_{g_{t}}^{(l)},^{4}W_{g_{t}}^{(l)}\end{bmatrix},b_{g_{t}}^{(l)}=\begin{bmatrix}{}^{1}b_{g_{t}}^{(l)}\\ {}^{2}b_{g_{t}}^{(l)}\end{bmatrix},$
where ${}^{1}W_{g_{t}}^{(l)},^{1}b_{g_{t}}^{(l)}$ have the same shapes as $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)}$ respectively, for $l=1,2,\ldots,l_{p}$, with the detailed initialization strategy as follows.
Input: OBS data $\mathcal{D}^{OBS}=\{(X_{i},T_{i},Y_{i},G_{i}=0)\}_{i=1}^{n}$, RCT data $\mathcal{D}^{RCT}=\{(X_{i},T_{i},Y_{i},G_{i}=1)\}_{i=n+1}^{n+m}$ and four hyperparameters $\lambda_{k}>0,k=1,\ldots,4.$
Compute $w_{i}=\frac{T_{i}}{2u}+\frac{1-T_{i}}{2(1-u)}$ with $u=\frac{1}{n}\sum_{i=1}^{n}T_{i}$ for $i=1,...,n$;
for number of steps for training the first-stage model do
      Sample a batch $\{(X_{i},T_{i},Y_{i})\}_{i\in B}$ from $\mathcal{D}^{OBS}$;
       Update $\theta_{1}=(\theta_{\phi},\theta_{\psi},\theta_{h_{0}},\theta_{h_{1}})$ by descending along the gradient $\nabla_{\theta_{1}}(\mathcal{L}_{f}+\lambda_{1}\mathcal{L}_{rec}+\lambda_{2}\mathcal{L}_{unb})$;
end for
Initialize $(^{1}W_{g_{t}}^{(l)},^{2}W_{g_{t}}^{(l)},^{3}W_{g_{t}}^{(l)},^{4}W_{g_{t}}^{(l)},^{1}b_{g_{t}}^{(l)},^{2}b_{g_{t}}^{(l)})$ by $(W_{h_{t}}^{(l)},0,0,0,b_{h_{t}}^{(l)},0)$ for $l=1,2,\ldots,l_{p}$ and $t=0,1$;
Compute $w_{i}=\frac{T_{i}}{2u}+\frac{1-T_{i}}{2(1-u)}$ with $u=\frac{1}{m}\sum_{i=n+1}^{n+m}T_{i}$ for $i=n+1,...,n+m$;
for number of steps for training the second-stage model do
      Sample a batch $\{(X_{i},T_{i},Y_{i})\}_{i\in B}$ from $\mathcal{D}^{RCT}$;
       Update $\theta_{2}=(\theta_{\phi^{U}},\theta_{g_{0}},\theta_{g_{1}})$ by descending along the gradient $\nabla_{\theta_{2}}(\mathcal{L}_{pred}+\lambda_{3}\mathcal{L}_{MI}+\lambda_{4}\mathcal{L}_{shift})$;
end for
Algorithm 1 Learning algorithm of the TSPF framework.
Initialization.
The initialization of the model parameters is crucial for the preservation of covariate information from the first stage as well as the effectiveness of the finetuning stage. The goal of initialization is to make sure the model initially produces the same prediction as the trained first-stage model. Nonetheless, a challenge of parameter initialization is that the model architecture of the second stage differs from that of the first stage, because of the augmented covariate representation. Based on the division of the parameters, we propose the following initialization strategy:
$(^{1}W_{g_{t}}^{(l)},^{2}W_{g_{t}}^{(l)},^{3}W_{g_{t}}^{(l)},^{4}W_{g_{t}}^{(l)},^{1}b_{g_{t}}^{(l)},^{2}b_{g_{t}}^{(l)})\leftarrow(W_{h_{t}}^{(l)},0,0,0,b_{h_{t}}^{(l)},0)$
for $l=1,2,\ldots,l_{p}$. That is, the shared parameters between the prediction heads $g_{t}$ and $h_{t}$ are initialized to be the same, and the rest parameters of the prediction head $g_{t}$ are initialized to be zero.
As in the first stage, we denote $f_{g}(x,t)=g_{t}([\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]^{\intercal})$ for $t\in\{0,1\}$, where $[\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]$ is the concatenated covariate representation. Given the RCT data $\{(X_{i},T_{i},Y_{i},G_{i}=1):i\in\mathcal{A}\}$, the prediction loss is computed as:
$\mathcal{L}_{pred}=\frac{1}{|\mathcal{R}|}\sum_{i\in\mathcal{R}}w_{i}\cdot l(Y_{i},f_{g}(X_{i},T_{i})),$
similarly with $w_{i}=\frac{T_{i}}{2u}+\frac{1-T_{i}}{2(1-u)}$ and $u=\frac{1}{m}\sum_{i=n+1}^{n+m}T_{i}$.
Regularization.
The presence of unmeasured confounding may cause a slight shift in the distribution of OBS data from the RCT data. Therefore the second-stage fine-tuned model should not deviate significantly from the first-stage model. We denote the initial value of $\theta_{g_{t}}$ as $\theta_{g_{t}}^{0}$. In order to constrain the deviation from the initial value, we include an $l_{2}-$norm in the loss function:
$\mathcal{L}_{shift}=||\theta_{g_{0}}-\theta_{g_{0}}^{0}||_{2}^{2}+||\theta_{g_{1}}-\theta_{g_{1}}^{0}||_{2}^{2}.$
Overall, the training objective of the second stage is given by:
$\min_{\phi^{U},~{}g_{0},~{}g_{1}}\mathcal{L}_{pred}+\lambda_{3}\mathcal{L}_{MI}+\lambda_{4}\mathcal{L}_{shift},$
where $\lambda_{3}>0$ and $\lambda_{4}>0$ are tunable hyperparameters. Note that during the second-phase training, we froze the parameters of the representation module $\phi$ and the decoder network $\psi$, while train the representation adapter module $\phi^{U}$ and the two prediction heads $g_{0},g_{1}$. We summarize the whole learning algorithm in Alg. 1.
Compared to residual correction methods as in Kallus et al. (Kallus et al., 2018), our representation adapter module guarantees a stronger representation ability, relaxing the linearly additive assumption. While compared to methods that jointly learn residual and CATE, one advantage is that when RCT data are limited, our proposed partial initialization strategy in the TSPF framework can avoid overfitting.
Table 1. The experiment results on the IHDP dataset and Jobs dataset. The best result is bolded and the second best is underlined.
IHDP Jobs
In-sample Out-sample In-sample Out-sample
Methods $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$
T-learner 0.44 $\pm$ 0.03 0.04 $\pm$ 0.02 0.52 $\pm$ 0.05 0.02 $\pm$ 0.01 0.66 $\pm$ 0.27 0.02 $\pm$ 0.02 0.60 $\pm$ 0.20 0.02 $\pm$ 0.01
S-learner 0.98 $\pm$ 0.18 0.04 $\pm$ 0.03 1.37 $\pm$ 0.34 0.15 $\pm$ 0.10 0.70 $\pm$ 0.30 0.02 $\pm$ 0.02 0.67 $\pm$ 0.39 0.03 $\pm$ 0.02
DR-learner 0.71 $\pm$ 0.19 0.06 $\pm$ 0.04 0.81 $\pm$ 0.25 0.06 $\pm$ 0.03 0.50 $\pm$ 0.07 0.07 $\pm$ 0.02 0.48 $\pm$ 0.09 0.07 $\pm$ 0.03
Causal Forest 1.90 $\pm$ 0.29 0.07 $\pm$ 0.05 2.04 $\pm$ 0.42 0.19 $\pm$ 0.11 1.38 $\pm$ 0.40 0.13 $\pm$ 0.08 1.23 $\pm$ 0.35 0.14 $\pm$ 0.08
TARNet 0.42 $\pm$ 0.09 0.05 $\pm$ 0.04 0.44 $\pm$ 0.12 0.05 $\pm$ 0.05 0.19 $\pm$ 0.17 0.02 $\pm$ 0.01 0.13 $\pm$ 0.02 0.02 $\pm$ 0.01
CEVAE 2.89 $\pm$ 0.72 0.15 $\pm$ 0.12 2.87 $\pm$ 0.80 0.22 $\pm$ 0.16 2.86 $\pm$ 0.80 0.35 $\pm$ 0.24 2.82 $\pm$ 0.74 0.46 $\pm$ 0.38
SCIGAN 2.53 $\pm$ 0.47 0.63 $\pm$ 0.29 2.58 $\pm$ 0.57 0.55 $\pm$ 0.48 2.28 $\pm$ 0.75 0.56 $\pm$ 0.15 2.15 $\pm$ 0.80 0.47 $\pm$ 0.21
DragonNet 0.19 $\pm$ 0.04 0.03 $\pm$ 0.02 0.26 $\pm$ 0.08 0.04 $\pm$ 0.02 0.15 $\pm$ 0.11 0.01 $\pm$ 0.01 0.11 $\pm$ 0.03 0.01 $\pm$ 0.01
DESCN 0.28 $\pm$ 0.06 0.05 $\pm$ 0.05 0.41 $\pm$ 0.11 0.07 $\pm$ 0.06 0.44 $\pm$ 0.09 0.28 $\pm$ 0.13 0.44 $\pm$ 0.08 0.27 $\pm$ 0.13
DRCFR 0.74 $\pm$ 0.32 0.15 $\pm$ 0.10 0.90 $\pm$ 0.52 0.18 $\pm$ 0.17 0.91 $\pm$ 0.49 0.08 $\pm$ 0.08 0.71 $\pm$ 0.43 0.09 $\pm$ 0.07
Twostep linear 0.64 $\pm$ 0.13 0.31 $\pm$ 0.23 0.87 $\pm$ 0.18 0.56 $\pm$ 0.24 0.69 $\pm$ 0.35 0.20 $\pm$ 0.12 0.56 $\pm$ 0.27 0.25 $\pm$ 0.16
CorNet 0.34 $\pm$ 0.12 0.05 $\pm$ 0.03 0.32 $\pm$ 0.09 0.04 $\pm$ 0.04 0.21 $\pm$ 0.09 0.04 $\pm$ 0.04 0.22 $\pm$ 0.08 0.05 $\pm$ 0.03
TSFP (ours) 0.13 $\pm$ 0.02 0.02 $\pm$ 0.02 0.16 $\pm$ 0.04 0.04 $\pm$ 0.02 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01 0.06 $\pm$ 0.01 0.01 $\pm$ 0.01
5. Experiment
5.1. Datasets
Following previous studies (Shalit et al., 2017; Louizos et al., 2017; Yoon et al., 2018), we conduct experiments on two publicly available datasets, namely IHDP (Hill, 2011) and Jobs (Shalit et al., 2017). The IHDP is a semi-synthetic dataset for causal effect estimation. The dataset is based on the Infant Health and Development Program, where the covariates are obtained by a randomized experiment investigating the effect of home visits by specialists on future cognitive scores. It consists of 747 units (19% treated, 81% control) and 25 covariates measuring the children and their mothers. The Jobs is a common benchmark dataset developed by LaLonde in 1986, studying the change of income and employment status after job training. We use an extended version of Jobs that comprises about 3,000 units (10% treated, 90% control) with 17 covariates.
5.2. Data Preprocessing
For both IHDP and Jobs, we simulate unmeasured confounding by generating a $c$-dimensional confounder $U_{i}\in\mathbb{R}^{c}$. To make sure the $U_{i}$ has a non-zero effect on $Y_{i}$ and $T_{i}$, we generate the data below:
$W_{1}\sim\mathcal{N}(0,0.1)^{d},~{}~{}~{}W_{2}\sim\mathcal{N}(0.02,0.1)^{c},~{}~{}~{}W_{3}\sim\mathcal{N}(0.1,1)^{d},$
$W_{4}\sim\mathcal{N}(0.1,1)^{c},~{}~{}~{}W_{5}\sim\mathcal{U}(0,0.2)^{d},~{}~{}~{}W_{6}\sim\mathcal{U}(0,0.2)^{c},$
$U_{i}\sim\mathcal{U}(0,0.2)^{c},~{}~{}~{}T_{i}\sim\text{Bern}(\sigma(W_{1}\cdot X_{i}+W_{2}\cdot U_{i})),$
$\mu_{i}^{0}=W_{3}\cdot X_{i}+W_{4}\cdot U_{i},~{}~{}~{}\mu_{i}^{1}=W_{5}\cdot X_{i}+W_{6}\cdot U_{i}+4,$
$Y_{i}^{0}\sim\mathcal{N}(\mu_{i}^{0},0.1),~{}~{}~{}Y_{i}^{1}\sim\mathcal{N}(\mu_{i}^{1},0.1),$
where $\mathcal{N}(\mu,D)$ denotes the normal distribution with mean $\mu$ and variance $D$, $\mathcal{U}(a,b)$ is the uniform distribution on interval $(a,b)$, $\text{Bern}(p)$ means the Bernoulli distribution with probability $p$, $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. Note that we keep $E[Y_{i}^{1}]=E[Y_{i}^{0}]+4$ as the same as the IHDP dataset and we let $E[W_{i}]>0\text{ for }i\in\{2,4,6\}$ to ensure the non-zero effect of $U_{i}$. The unmeasured confounding strength parameter $c$ is set to 30. Then we slice the training, validation, and test sets in the ratio of 63/27/10. In addition, to obtain a separate RCT training dataset for data fusion, we first randomly split $r\%$ of the training samples, and then assign treatments $T^{new}_{i}$ according to the following formula and replace the factual treatment $T_{i}$ and outcome $Y^{f}_{i}$ to obtain a RCT dataset:
$T^{new}_{i}=\operatorname{Bern}(0.5),~{}~{}~{}Y^{new}_{i}=\mathbb{I}(T^{new}_{i}=T_{i})(Y^{f}_{i}-Y^{cf}_{i})+Y^{cf}_{i},$
where $\mathbb{I}(\cdot)$ is the indicator function, $Y^{f}_{i}=T_{i}Y_{i}^{1}+(1-T_{i})Y_{i}^{0}$ is the factual outcome, and $Y^{cf}_{i}=T_{i}Y_{i}^{0}+(1-T_{i})Y_{i}^{1}$ is the counterfactual outcome. The RCT data ratio $r\%$ is set to 10% unless otherwise stated. Finally, we replace the $T_{i}$ and $Y_{i}$ using the above formula for all samples in the validation set.
Table 2. Ablation study on the pretraining stage and fine-tuning stage, the initialization strategy, and the training strategy of freezing foundational representation module and the initialized parameters.
IHDP Jobs
In-sample Out-sample In-sample Out-sample
$\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$
TSFP w/o $\mathcal{L}_{rec}$ and $\mathcal{L}_{imb}$ 0.25 $\pm$ 0.05 0.07 $\pm$ 0.02 0.29 $\pm$ 0.08 0.07 $\pm$ 0.02 0.19 $\pm$ 0.03 0.04 $\pm$ 0.03 0.18 $\pm$ 0.02 0.04 $\pm$ 0.03
TSFP w/o $\mathcal{L}_{rec}$ $0.21\pm 0.04$ $0.05\pm 0.02$ $0.24\pm 0.05$ $0.05\pm 0.02$ $0.14\pm 0.01$ $0.03\pm 0.01$ $0.13\pm 0.02$ $0.03\pm 0.01$
TSFP w/o $\mathcal{L}_{imb}$ $0.18\pm 0.03$ $0.03\pm 0.02$ $0.19\pm 0.05$ $0.05\pm 0.02$ $0.12\pm 0.02$ $0.03\pm 0.02$ $0.10\pm 0.04$ $0.03\pm 0.01$
TSFP w/o $\mathcal{L}_{MI}$ and $\mathcal{L}_{shift}$ 0.23 $\pm$ 0.03 0.06 $\pm$ 0.01 0.26 $\pm$ 0.05 0.07 $\pm$ 0.02 0.16 $\pm$ 0.04 0.03 $\pm$ 0.01 0.18 $\pm$ 0.04 0.03 $\pm$ 0.01
TSFP w/o $\mathcal{L}_{MI}$ 0.17 $\pm$ 0.04 0.04 $\pm$ 0.01 0.18 $\pm$ 0.04 0.05 $\pm$ 0.02 0.12 $\pm$ 0.03 0.02 $\pm$ 0.01 0.15 $\pm$ 0.03 0.02 $\pm$ 0.01
TSFP w/o $\mathcal{L}_{shift}$ 0.19 $\pm$ 0.04 0.05 $\pm$ 0.01 0.21 $\pm$ 0.05 0.05 $\pm$ 0.02 0.15 $\pm$ 0.05 0.03 $\pm$ 0.01 0.16 $\pm$ 0.04 0.02 $\pm$ 0.01
TSFP not freezing representation $0.17\pm 0.02$ $0.03\pm 0.03$ $0.20\pm 0.05$ $0.05\pm 0.04$ 0.10 $\pm$ 0.03 0.01 $\pm$ 0.01 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01
TSFP freezing initialized parameters $0.16\pm 0.08$ $0.03\pm 0.02$ $0.19\pm 0.05$ $0.04\pm 0.02$ 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01 0.08 $\pm$ 0.04 0.01 $\pm$ 0.01
TSFP w/o initialization $0.35\pm 0.09$ $0.08\pm 0.07$ $0.38\pm 0.11$ $0.07\pm 0.06$ 0.42 $\pm$ 0.12 0.06 $\pm$ 0.04 0.41 $\pm$ 0.13 0.06 $\pm$ 0.04
TSFP w/o initialization and $\mathcal{L}_{shift}$ $0.43\pm 0.09$ $0.11\pm 0.07$ $0.45\pm 0.10$ $0.12\pm 0.06$ 0.54 $\pm$ 0.13 0.09 $\pm$ 0.05 0.53 $\pm$ 0.14 0.11 $\pm$ 0.05
TSFP w/o $Z^{U}$ $0.18\pm 0.03$ $0.05\pm 0.02$ $0.19\pm 0.04$ $0.05\pm 0.02$ 0.14 $\pm$ 0.03 0.02 $\pm$ 0.01 0.16 $\pm$ 0.04 0.02 $\pm$ 0.01
TSFP $0.13\pm 0.02$ $0.02\pm 0.02$ $0.16\pm 0.04$ $0.04\pm 0.02$ 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01 0.06 $\pm$ 0.01 0.01 $\pm$ 0.01
Figure 2. The experiment results for the IHDP dataset under different RCT data ratios in the training data.
5.3. Baselines and Evaluation Metrics
5.3.1. Baselines
  • T-learner (Künzel et al., 2019): T-learner utilizes two separate regressors for each treatment group.
  • S-learner (Athey and Imbens, 2015): S-learner treats the indicator of treatment as features, utilizing a single model to estimate the potential outcome for both treatment and control groups.
  • DR-learner (Kennedy, 2023): DR-learner estimates the CATE via cross-fitting a doubly robust score function in two stages.
  • SCIGAN (Yoon et al., 2018): SCIGAN utilizes a generative adversarial network to model treatment effect.
  • Causal Forest (Wager and Athey, 2018): Causal Forest is a random forest-based model that directly estimates the treatment effect.
  • TARNet (Shalit et al., 2017): TARNet applies a shared representation layer and a two-head network inference layer.
  • DragonNet (Shi et al., 2019): DragonNet designs an adaptive neural network to learn propensities and counterfactual outcomes.
  • DESCN (Zhong et al., 2022): DESCN uses deep networks to model treatment effects in the entire sample space.
  • DRCFR (Hassanpour and Greiner, 2019): DRCFR aims to learn disentangled representations and address selection bias in CATE estimation.
  • Twostep linear (Kallus et al., 2018): Twostep linear method uses OBS data to learn a biased estimate for the treatment effect and then aims to remove the bias using RCT data.
  • CorNet (Hatt et al., 2022b): CorNet uses the RCT data to learn a non-linear bias function in the second stage.
5.3.2. Evaluation Metrics
Following previous studies (Shalit et al., 2017; Yao et al., 2018), we evaluate the performance of CATE estimation using the square root of Precision in Estimation of Heterogeneous Effects (PEHE):
$\sqrt{{\epsilon_{\text{PEHE}}}}=\sqrt{{\frac{1}{n}\sum_{i=1}^{n}((\hat{Y}_{i}^{1}-\hat{Y}_{i}^{0})-(Y_{i}^{1}-Y_{i}^{0}))^{2}}},$
where $\hat{Y}_{i}^{t}$ and $Y_{i}^{t}$ are the predicted and ground truth values for the potential outcomes of individual $i$ under treatment $t$. In addition, we also use the absolute error in Average Treatment Effect (ATE) to evaluate estimation performance, which is defined as:
$\epsilon_{\text{ATE}}=\frac{1}{n}\left|\sum_{i=1}^{n}((\hat{Y}_{i}^{1}-\hat{Y}_{i}^{0})-(Y_{i}^{1}-Y_{i}^{0}))\right|.$
The smaller the evaluation metrics, the better the estimation.
5.3.3. Experimental Details
We implement TSPF with a multi-layer perceptron (Haykin, 1994) with 2 layers for our representation and reconstruction modules as well as the prediction heads in both stages using the PyTorch framework. We tune the hyper-parameters in the loss functions from $51e-$ 5 to $0.1$, the learning rate from $51e-$ 5 to $21e-$ 2, and the weight decay parameter from $51e-$ 5 to $21e-$ 2. The batch size is set to 256 in all scenarios and the optimizer is selected in Adam and SGD. We report both in-sample and out-of-sample results for metrics $\sqrt{\epsilon_{\mathrm{PEHE}}}$ and $\epsilon_{\mathrm{ATE}}$ in our experiments.
5.4. Performance Analysis
Table 1 shows the prediction performance with varying baselines and our methods. First, representation-based methods generally outperform generation-based methods and meta-learners, which shows the effectiveness of causal representation learning. Note that our proposed TSPF exhibits the most competitive performance in most cases, outperforming all the baselines except for the $\epsilon_{\mathrm{ATE}}$ metric in out-sample scenario. In the comparison of two-step methods, we can see that CorNet (Hatt et al., 2022b), which adopts a nonlinear residual module, outperforms the Twostep linear method based on the assumption that the residual $\eta(X_{i})$ is linear and additive. In addition, our method stably outperforms the other two-stage methods, showing the effectiveness of our pretraining-finetuning framework.
5.5. Ablation Study
After evaluating the overall performance of our method, we perform ablation studies to validate the effectiveness of each module in our approach. Firstly, we consider employing all the modules in the first stage and taking away the mutual information term and/or the shift regularization term. Then we include all the modules in the second stage and test the effect of removing the reconstruction module and/or the IPM distance. To validate the soundness of our training algorithm, especially the initialization strategy, we also evaluate the performance of TSPF when not freezing the foundational representation module $\phi$, freezing the initialized parameters in $g_{0}$ and $g_{1}$, without the initialization strategy, and without the augmented representation $Z^{U}$. The results are presented in Table 2, which shows that each component in our framework plays an important role in accurately estimating CATE. In the first/second stage, $\mathcal{L}_{rec}$/$\mathcal{L}_{shift}$ is more important than $\mathcal{L}_{imb}$/$\mathcal{L}_{MI}$. In addition, our initialization strategy significantly improves the performance.
5.6. In-Depth Analysis
Table 3. Experimental results on the IHDP dataset with unmeasured confounding strength $c=10$ and $c=50$. The best result is bolded and the second best is underlined.
unmeasured confounding strength $c=10$ unmeasured confounding strength ${c=50}$
In-sample Out-sample In-sample Out-sample
Methods $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$
T-learner 0.42 $\pm$ 0.06 0.03 $\pm$ 0.02 0.47 $\pm$ 0.08 0.06 $\pm$ 0.05 0.48 $\pm$ 0.05 0.05 $\pm$ 0.02 0.53 $\pm$ 0.06 0.04 $\pm$ 0.03
S-learner 0.99 $\pm$ 0.26 0.05 $\pm$ 0.04 1.29 $\pm$ 0.34 0.13 $\pm$ 0.11 1.17 $\pm$ 0.11 0.05 $\pm$ 0.03 1.51 $\pm$ 0.19 0.11 $\pm$ 0.10
DR-learner 0.69 $\pm$ 0.12 0.06 $\pm$ 0.03 0.72 $\pm$ 0.13 0.10 $\pm$ 0.05 0.73 $\pm$ 0.11 0.07 $\pm$ 0.04 0.80 $\pm$ 0.13 0.08 $\pm$ 0.06
Causal Forest 1.93 $\pm$ 0.31 0.08 $\pm$ 0.07 1.85 $\pm$ 0.28 0.24 $\pm$ 0.22 1.98 $\pm$ 0.18 0.13 $\pm$ 0.06 2.04 $\pm$ 0.34 0.29 $\pm$ 0.19
TARNet 0.26 $\pm$ 0.06 0.03 $\pm$ 0.02 0.27 $\pm$ 0.13 0.04 $\pm$ 0.04 0.52 $\pm$ 0.20 0.09 $\pm$ 0.09 0.53 $\pm$ 0.24 0.11 $\pm$ 0.11
CEVAE 3.09 $\pm$ 0.92 0.19 $\pm$ 0.16 3.19 $\pm$ 1.22 0.41 $\pm$ 0.33 3.50 $\pm$ 0.51 0.25 $\pm$ 0.16 3.39 $\pm$ 0.45 0.65 $\pm$ 0.32
SCIGAN 2.74 $\pm$ 0.68 0.62 $\pm$ 0.28 2.77 $\pm$ 0.89 0.56 $\pm$ 0.30 2.94 $\pm$ 0.42 0.46 $\pm$ 0.25 2.90 $\pm$ 0.35 0.75 $\pm$ 0.38
DragonNet 0.19 $\pm$ 0.04 0.03 $\pm$ 0.02 0.23 $\pm$ 0.05 0.06 $\pm$ 0.07 0.23 $\pm$ 0.05 0.06 $\pm$ 0.07 0.29 $\pm$ 0.13 0.07 $\pm$ 0.04
DESCN 0.36 $\pm$ 0.17 0.04 $\pm$ 0.03 0.42 $\pm$ 0.26 0.08 $\pm$ 0.09 0.46 $\pm$ 0.14 0.09 $\pm$ 0.05 0.55 $\pm$ 0.23 0.08 $\pm$ 0.06
DRCFR 0.70 $\pm$ 0.32 0.09 $\pm$ 0.08 0.92 $\pm$ 0.74 0.18 $\pm$ 0.22 1.21 $\pm$ 0.80 0.30 $\pm$ 0.20 1.17 $\pm$ 0.72 0.26 $\pm$ 0.18
Twostep linear 0.70 $\pm$ 0.16 0.29 $\pm$ 0.25 0.83 $\pm$ 0.30 0.30 $\pm$ 0.39 0.76 $\pm$ 0.21 0.42 $\pm$ 0.23 0.83 $\pm$ 0.25 0.33 $\pm$ 0.31
CorNet 0.23 $\pm$ 0.12 0.05 $\pm$ 0.04 0.27 $\pm$ 0.13 0.04 $\pm$ 0.04 0.39 $\pm$ 0.19 0.08 $\pm$ 0.08 0.38 $\pm$ 0.18 0.07 $\pm$ 0.07
TSFP (ours) 0.15 $\pm$ 0.04 0.02 $\pm$ 0.02 0.15 $\pm$ 0.04 0.02 $\pm$ 0.02 0.15 $\pm$ 0.03 0.02 $\pm$ 0.01 0.17 $\pm$ 0.07 0.02 $\pm$ 0.02
Figure 3. Sensitivity analysis on the reconstruction loss strength $\lambda_{1}$, covariate imbalancing loss strength $\lambda_{2}$, mutual information loss strength $\lambda_{3}$ and RCT shift loss strength $\lambda_{4}$ on the IHDP dataset.
5.6.1. RCT Data Ratio
To validate whether our framework performs well with limited RCT data, we further explore the effect of using different ratios of RCT data in the training set. We conduct experiments with a sequence of increasing ratios from 2% to 20%. For each scenario, the optimal network size and hyper-parameters are finetuned to ensure fair comparison. The results in Figure 2 show that our method is robust to the change of RCT data ratio and outperforms the baselines under different RCT ratios.
5.6.2. Unmeasured Confounding Strength
Our approach aims to achieve unbiased estimation of CATE in the presence of unmeasured confounding. To verify the validity under different strengths of unmeasured confounding, we add a comparison of the performance under low ($c=10$) and high ($c=50$) unmeasured confounding effect, which is shown in Table 3. The results show that our method stably outperforms baselines with varying $c$.
5.7. Sensitivity Analysis
Despite there are many hyper-parameters in our TSPF framework, they present in two separate stages, i.e, $\lambda_{1}$ and $\lambda_{2}$ in the first stage, while $\lambda_{3}$ and $\lambda_{4}$ in the second stage. As a result, the hyper-parameter search space is significantly reduced, compared with simultaneously searching four hyper-parameters. We explore the sensitivity of the hyper-parameters on the IHDP dataset and the results are shown in Figure 3. For all four hyper-parameters, the best result is achieved with moderate values. The performance dramatically drops with low values of hyper-parameters, which further shows the effectiveness of each component in our method.
6. Conclusion
This paper studies the CATE estimation in the presence of unmeasured confounding fusing both large-scale OBS data and small-scale RCT data. We propose a two-stage pretraining-finetuning framework to tackle the overfitting problem caused by the small-scale RCT data. Specifically, the foundational representation learned in the first stage is used to adjust for the measured confounding in the OBS data. The augmented representation learned in the second stage is used to adjust for the unmeasured confounding guided by the RCT data. To avoid overfitting caused by the small-scale RCT data in the second stage, instead of training a separate network, we further propose to partially initialize the network parameters from the pretrained network from the first stage. Compared to the previous CATE estimation methods that combine OBS and RCT data, our approach has the advantage of not restricting the data-generating process (e.g., linearity or additive noise assumptions) and not suffering from overfitting. The extensive semi-synthetic and real-world experiments conducted on two widely-used public datasets demonstrate the superiority of our method.
Acknowledgements.
HL was supported by National Natural Science Foundation of China (623B2002). MG was supported by ARC DE210101624 and ARC DP240102088.